Can you help me find the ordered pair by using the substitution method
5x+4y= -2
x=29-5y

Answer by rfer(16322)

5x+4y=-2
x=29-5y
----------
5(29-5y)+4y=-2
145-25y+4y=-2
145-21y=-2
147=21y
y=7
--------------
x=29-5(7)
x=29-35
x=-6
(-6,7)